1. A lamp (R = 147 ohms), an electric heater (R = 24 ohms), and a fan (R = 49 ohms) are connected in parallel to an extension cord plugged into a 117 V outlet.

(a) What total current is supplied to the circuit?

(b) What is the voltage across the fan?

(c) What is the current in the lamp?

(d) What power is expended in the heater?
Answer #1

As al three are in Parallel so 1/R​​​​​​eq = (1/R​​​​​​lamp) + (1/R​​​​​​heater) + (1/R​​​​​​fan)

=> 1/R​​​​​​eq = (1/147) + (1/24) + (1/49)

=> R​​​​​eq = 14.5 ohms

a. I = V/R​​​​​​eq = 117 V / 14.5 ohms = 8 A

b. If all are parallel so Voltage across all would be same as total voltage. So, V​​​​​​fan = 117 V

c. I​​​​​​fan = V/R​​​​​​fan = 117 V / 49 ohms = 2.39 A

d. Power​​heater= V²/R​​​​​​heater = (117 V)²/ 24 ohms = 570.37 W

If you find this Solution helpful then please give a thumbsup ?. Thankyou.